OpenStudy (anonymous):

Solve, (x-1) / (x+1) > (x + 3) / (2x + 3)

OpenStudy (anonymous):

(x-1)(2x+3)-(x=3)(x+1)>0

OpenStudy (anonymous):

(x-1)/(x+1) - (x+3)/(2x+3) > 0 (x-1)(2x+3) - (x+3)(x+1) > 0\[(-\infty,(3-\sqrt{33})/2 )\cup((3+\sqrt{33})/2 ,\infty)\]

OpenStudy (anonymous):

(x-1)(2x+3)-(x+3)(x+1)>0 x^2-3x-6>0 x=(3+sqrt{33})/2 or =(3-sqrt{33})/2

OpenStudy (anonymous):

(x-1) / (x+1) > (x + 3) / (2x + 3)

OpenStudy (anonymous):

take cross mutiply

OpenStudy (anonymous):

(x-1)(2x+3)>(x+3)(x+1)

OpenStudy (anonymous):

(2x^2+x-3) > (x^2 +4x +3)

OpenStudy (anonymous):

(2x^2+x-3) - (x^2 +4x +3)>0

OpenStudy (anonymous):

x^2 -3x -6>0

OpenStudy (anonymous):

using quadratic formula

OpenStudy (anonymous):

x = (-b +-(sqrt(b^2-4ac)))/2a

OpenStudy (anonymous):

Answer is x > 1/2( 3 +sqrt(33)

OpenStudy (anonymous):

x < 1/2(3 - sqrt(33))